I need help with these two problems.
1. 10m^4n^2-2m^3n^3-4m^2n^4
2. Find three consecutive odd integers such that three times the first minus the second is one more than the third.

Answer by checkley77(12844)

1. 10m^4n^2-2m^3n^3-4m^2n^4
2M^2N^2(5M^2-MN-2N^2) ANS.
2. Find three consecutive odd integers such that three times the first minus the second is one more than the third.
LET X, X+2 & X+4 BE THE THREE CONSECUTIVE ODD INTEGERS.
3X-(X+2)=(X+4)-1
3X-X-2=X+3
2X-X=3+2
X=5 ANS. FOR THE FIRST ODD INTEGER.
5+2=7 FOR THE SECOND
5+4=9 FOR THE THIRD.
PROOF:
3*5-7=9-1
15-7=8
8=8